# MACRO-LLM: LLM-Empowered Multi-Agent Collaborative Reasoning under Spatiotemporal Partial Observability

### v1

@@ Appendix B Detailed Algorithm of Proposal Generation
The computational complexity of the CoProposer is dominated by the number of LLM inference calls. Let $Att_{\text{max}}$ be the maximum number of validation attempts and $K$ be the rollout horizon. In the worst-case scenario (where valid proposals are hard to find), the complexity is $\mathcal{O}(Att_{\text{max}}\cdot k)$. To mitigate latency in real-world applications, we set $Att_{\text{max}}=10$ for all task. The binary simplification of rewards beyond $t+1$ further reduces the overhead of reward calculation.
@@ 17      Add $(a_{n}^{t+k,att},o_{n}^{t+k,att},R_{n}^{t+k,att})$ to rollout trajectory $\tau_{n}^{t,att}$;
Appendix C Equations for Welford-based State Aggregation
@@
Appendix D Simulation Environments and Task Configurations
We evaluate our framework on two distinct and challenging domains to assess its collaborative performance and generalization capabilities, utilizing high-fidelity simulators for each task.
D.1 Cooperative Adaptive Cruise Control (CACC)
The CACC task is simulated using the Simulation of Urban Mobility (SUMO) [^1] SUMO: https://eclipse.dev/sumo/, licensed under EPL 2.0. (Version 1.21.0).
  • Topology: We model a platoon of eight vehicles ($N=8$) on a single-lane highway segment. The simulation runs for a duration of 60s with a time step of 0.5s (i.e., 120 time steps in total). Each agent operates with limited perception, accessing only the state information (e.g., position, velocity) of its immediately adjacent vehicles (predecessor and follower).
  • Objective: The control objective is for agents to autonomously adjust acceleration to stabilize the platoon, converging rapidly to an optimal headway of 20 m and a target velocity of 15 m/s.
@@ D.2 Pandemic Control (PC)
  • Topology: We model the urban environment as a graph of seven agents representing distinct facility types: Home, Office, School, Hospital, Retail, Restaurant, and Government. To construct realistic urban topologies, we utilize Google Maps to identify clusters of functionally similar buildings within specific urban blocks, marking their centroids as individual agents. Communication links are established based on spatial proximity, forming a connected topology as shown in Fig. 6. The simulation spans 120 days, during which agents operate under partial observability regarding global infection rates.
  • Objective: The primary objective is for agents to dynamically adjust regulatory policies (e.g., gathering restrictions, business hours) to minimize the spread of the virus while maintaining essential socio-economic activities.
@@
Appendix E Baseline Methods
To evaluate MACRO-LLM, we implement and adapt the following baselines for both the CACC and PC tasks.
E.1 MARL Baselines
@@ E.2 LLM-based MAS Baselines
We compare against recent frameworks utilizing LLMs for collaboration: (1) ToM-Belief Li et al. (2023b): A MAS that employs second-order Theory of Mind for introspection via belief state updates and environmental feedback; (2) ChatEval Chan et al. (2023): Adopts the Simultaneous-Talk-with-Summarizer protocol (Algorithm 3), which is structurally similar to our task setting; (3) LAMEN Davidson et al. (2024): A framework that constructs internal "mental notes" alongside public messages to facilitate strategic negotiation.
E.3 Adaptation and Fairness
To ensure a fair comparison, we maintain experimental consistency across all LLM-based methods. Specifically, we use standardized core prompts for equivalent procedural steps while integrating unique prompting techniques specific to each baseline (e.g., structures for mental notes generation in LAMEN or belief updates in ToM-Belief). All baselines utilize the same backbone model (GPT-4o) and environmental interfaces as MACRO-LLM.
Appendix F Detailed Evaluation Metrics
@@ F.2 PC Task Metrics
Appendix G Computational and Communication Scalability Analysis
Figure 7: Training and execution time on the pandemic control task (120-day simulation), considering the total time from training to execution for a single node in a new environment. Experiment settings can be found in Sec. 4.2.
43+4
@@ Appendix G Computational and Communication Scalability Analysis
We provide a theoretical and empirical analysis of MACRO-LLM, focusing on communication complexity and computational workload.
G.1 Communication Complexity Analysis
@@
Following the definition by Yao Yao (1979), communication complexity is quantified by the total number of bits exchanged during the protocol. In our decentralized framework, we focus on the per-agent communication load. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be the agent communication graph. We denote the average node degree as $\bar{d}$ and the maximum negotiation rounds as $r$. The topology of the CACC task is linear ($\bar{d}\approx 2$). In the PC task, the topology represents an urban graph with average degree $\bar{d}_{\text{PC}}$. We assume sparse topologies where $\bar{d}$ does not grow linearly with the total number of agents $N$ (i.e., $\bar{d}\ll N$). As the network scales up, MACRO-LLM employs mean-field approximation to compress the states of unobservable agents into a single virtual node representation. Neglecting inference stochasticity, the per-agent message size is theoretically invariant to the network scale, which represented as $M$. Consequently, the per-agent communication complexity can be represented as $O(\bar{d}\cdot r\cdot M)$. Since $\bar{d}$, $r$, and $M$ are independent of network scale $N$, the per-agent load remains constant ($O(1)$) as the network scales in theory.
G.2 Computational Efficiency and Latency
@@
Appendix H Extended Ablation Analysis
@@
H.1 Component Analysis in CACC
The removal of specific modules leads to critical performance degradation in the high-precision and highly dynamic CACC task:
@@ H.1 Component Analysis in CACC
H.2 Component Analysis in PC
A similar pattern holds for the Pandemic Control task:
@@ H.2 Component Analysis in PC
Appendix I Analysis of Failure Cases and Solutions

### v2

@@ 17      Add $(a_{n}^{t+k,att},o_{n}^{t+k,att},R_{n}^{t+k,att})$ to rollout trajectory $\tau_{n}^{t,att}$;
The computational complexity of the CoProposer is dominated by the number of LLM inference calls. Let $Att_{\text{max}}$ be the maximum number of validation attempts and $K$ be the rollout horizon. In the worst-case scenario (where valid proposals are hard to find), the complexity is $\mathcal{O}(Att_{\text{max}}\cdot k)$. To mitigate latency in real-world applications, we set $Att_{\text{max}}=10$ for all tasks. The binary simplification of rewards beyond $t+1$ further reduces the overhead of reward calculation.
Appendix C Scope of Theoretical Assumptions
Eq. 4 analyzes the mean-field approximation error under the domain’s true reward function $R_{n}$, i.e., the error introduced by substituting the true neighborhood states $s^{t}_{\mathcal{N}_{n}}$ with the mean-field estimate $\mu_{n}$. We clarify the scope and assumptions below.
Domain Satisfaction.
In our two evaluation domains, the domain reward functions are governed by continuously differentiable physical dynamics: kinematic equations relating headway and velocity deviations in CPP, and epidemiological SIR-family dynamics governing infection rates in PC. Both operate over bounded state-action spaces with finite horizons, satisfying the smoothness requirement for the Taylor expansion argument. The constant $L_{2}$ is therefore determined by the continuity properties of these domain-specific dynamics.
Scope Clarification.
Eq. 4 characterizes the theoretical relationship between neighborhood variance $\sigma^{2}_{n}$ and the mean-field approximation quality under the domain’s true reward function. This analysis establishes $\sigma^{2}_{n}$ as a principled uncertainty indicator, motivating the framework’s design choice of transmitting $\sigma^{2}_{n}$ during negotiation. In practice, the LLM’s rollout-based reward estimation introduces a separate source of approximation that is outside the scope of Eq. 4. This estimation gap is mitigated through progressive constraint relaxation in rollout verification (Sec. 3.2), which simplifies reward signals beyond $t+1$ to binary feasibility indicators, reducing sensitivity to prediction imprecision.
Appendix D Proof of Theorem 1
Proof.
When the neighborhood $\mathcal{N}_{n}$ is separated into $P$ disjoint subgroups $\{\mathcal{N}_{n}^{(p)}\}_{p=1}^{P}$ and apply Eq. 4 independently within each subgroup. For each subgroup $p$, define the local weighted mean $\mu_{n}^{(p)}=\sum_{m\in\mathcal{N}_{n}^{(p)}}w_{n,m}^{(p)}\,s_{m}^{t}$ and local weighted variance $\sigma_{n}^{2,(p)}=\sum_{m\in\mathcal{N}_{n}^{(p)}}w_{n,m}^{(p)}\,(s_{m}^{t}-\mu_{n}^{(p)})^{2}$, where $w_{n,m}^{(p)}=w_{n,m}/W^{(p)}$ are the renormalized weights within subgroup $p$ and $W^{(p)}=\sum_{m\in\mathcal{N}_{n}^{(p)}}w_{n,m}$. By the same argument as in Eq. 4 (Taylor expansion with first-order cancellation within each subgroup), the contribution of subgroup $p$ to the approximation error is bounded by:
$\begin{split}|\Delta\mathcal{R}_{n}^{(p)}|&\leq\frac{L_{2}}{2}\sum_{m\in\mathcal{N}_{n}^{(p)}}(s_{m}^{t}-\mu_{n}^{(p)})^{2}\\ &\leq\frac{L_{2}}{2\,w_{\min}}\sum_{m\in\mathcal{N}_{n}^{(p)}}w_{n,m}\,(s_{m}^{t}-\mu_{n}^{(p)})^{2}.\end{split}$ (6)
Since the subgroups are disjoint, the total approximation error is bounded by:
$\begin{split}&\left|\mathcal{R}_{n}(\mathbf{s}_{\mathcal{N}_{n}}^{t})-\mathcal{R}_{n}(\boldsymbol{\mu}_{n}^{(1:P)})\right|\leq\sum_{p=1}^{P}|\Delta\mathcal{R}_{n}^{(p)}|\\ \leq&\frac{L_{2}}{2\,w_{\min}}\sum_{p=1}^{P}\sum_{m\in\mathcal{N}_{n}^{(p)}}w_{n,m}\,(s_{m}^{t}-\mu_{n}^{(p)})^{2}.\end{split}$ (7)
By the law of total variance Casella and Berger (2024), for any partition of a weighted dataset into $P$ groups:
$\begin{split}&\underbrace{\sum_{m\in\mathcal{N}_{n}}w_{n,m}\,(s_{m}^{t}-\mu_{n})^{2}}_{\sigma_{n}^{2}\text{ (total variance)}}\\ =&\underbrace{\sum_{p=1}^{P}W^{(p)}\sigma_{n}^{2,(p)}}_{\text{within-group variance}}+\underbrace{\sum_{p=1}^{P}W^{(p)}(\mu_{n}^{(p)}-\mu_{n})^{2}}_{\text{between-group variance}\;\geq\;0},\end{split}$ (8)
where $W^{(p)}=\sum_{m\in\mathcal{N}_{n}^{(p)}}w_{n,m}$. Since the between-group variance is non-negative:
$\sum_{p=1}^{P}W^{(p)}\sigma_{n}^{2,(p)}\leq\sigma_{n}^{2}.$ (9)
Furthermore, noting that $\sum_{m\in\mathcal{N}_{n}^{(p)}}w_{n,m}\,(s_{m}^{t}-\mu_{n}^{(p)})^{2}=W^{(p)}\sigma_{n}^{2,(p)}$, we can obtain:
$\begin{split}&\sum_{p=1}^{P}\sum_{m\in\mathcal{N}_{n}^{(p)}}w_{n,m}\,(s_{m}^{t}-\mu_{n}^{(p)})^{2}\\ =&\sum_{p=1}^{P}W^{(p)}\sigma_{n}^{2,(p)}\leq\sigma_{n}^{2}.\end{split}$ (10)
Substituting back into (7):
$\begin{split}&\left|\mathcal{R}_{n}(\mathbf{s}_{\mathcal{N}_{n}}^{t})-\mathcal{R}_{n}(\boldsymbol{\mu}_{n}^{(1:P)})\right|\\ \leq&\frac{L_{2}}{2\,w_{\min}}\sum_{p=1}^{P}W^{(p)}\sigma_{n}^{2,(p)}\leq\frac{L_{2}}{2\,w_{\min}}\cdot\sigma_{n}^{2}\\ =&\frac{L_{2}}{2\,w_{\min}}\cdot|\mathcal{N}_{n}|\cdot\bar{\sigma}_{n}^{2},\end{split}$ (11)
where $\bar{\sigma}_{n}^{2}=\sigma_{n}^{2}/|\mathcal{N}_{n}|$ is the per-neighbor average variance. The inequality is strict whenever the subgroup means $\mu_{n}^{(p)}$ differ, since the between-group variance $\sum_{p}W^{(p)}(\mu_{n}^{(p)}-\mu_{n})^{2}>0$ in that case. Therefore, if the topology can be divided into several disjoint group, the error bound is tighter. ∎
Appendix E Proof of Theorem 2 (Properties of the Negotiation Loop)
Proof.
$\mathcal{G}=(\mathcal{V},\mathcal{E})$ represents a connected communication graph, and $r$ denotes the maximum number of negotiation rounds.
Bounded Termination.
At each round $\ell$, the negotiation loop evaluates one of two termination conditions:
  • Early stopping via local consensus: If $\|P_{n}^{(\ell)}-P_{m}^{(\ell)}\|<\delta$ for all $m\in\mathcal{N}_{n}$, where $\delta$ is a predefined task-specific threshold (e.g., $\delta=0.1\,\text{m/s}^{2}$ for CPP), the agent terminates negotiation.
  • Maximum round limit: The loop terminates after at most $r$ rounds regardless of consensus status.
Since both conditions are finitely bounded, termination is guaranteed for every agent $n\in\mathcal{V}$.
Monotonic Non-Decrease of Proposal Quality.
At the beginning of each negotiation round $\ell+1$, the best proposal from the previous round, $P_{n}^{(\ell)}$, is retained in the candidate set for the current round. That is, the candidate set at round $\ell+1$ is:
$\mathcal{C}_{n}^{(\ell+1)}=\{P_{n}^{(\ell)}\}\cup\mathcal{C}_{n}^{\text{new}},$ (12)
where $\mathcal{C}_{n}^{\text{new}}$ denotes the set of newly generated proposals incorporating neighbor information received at round $\ell+1$.
Since $P_{n}^{(\ell)}\in\mathcal{C}_{n}^{(\ell+1)}$ by the Design Invariant, and agent $n$ selects $P_{n}^{(\ell+1)}=\arg\max_{P\in\mathcal{C}_{n}^{(\ell+1)}}\mathcal{R}(\tau_{n}(P))$, we have $\mathcal{R}(\tau_{n}^{(\ell+1)})=\max_{P\in\mathcal{C}_{n}^{(\ell+1)}}\mathcal{R}(\tau_{n}(P))\geq\mathcal{R}(\tau_{n}(P_{n}^{(\ell)}))\mathcal{R}(\tau_{n}^{(\ell)}).$ Therefore, the sequence $\{\mathcal{R}(\tau_{n}^{(\ell)})\}_{\ell=0}^{r}$ is monotonically non-decreasing.
The cumulative rollout reward is upper-bounded by the maximum achievable return over the rollout horizon $k$:
$\mathcal{R}(\tau_{n}^{(\ell)})\leq\sum_{\iota=0}^{k}\gamma^{\iota}R_{\max}=R_{\max}\cdot\frac{1-\gamma^{k+1}}{1-\gamma},$ (13)
where $R_{\max}=\sup_{o_{n},a_{n}}|R_{n}(o_{n},a_{n})|$ is the single-step reward bound. Therefore, $\{\mathcal{R}(\tau_{n}^{(\ell)})\}_{\ell=0}^{r}$ is a bounded, monotonically non-decreasing sequence of real numbers, which necessarily converges. ∎
Appendix F Equations for Welford-based State Aggregation
@@
Appendix G Simulation Environments and Task Configurations
Table 3 summarizes the key settings of the MACRO-LLM framework used across all experiments.
Table 3: Framework hyperparameters.
Parameter Value
Rollout horizon $k$ 2
Max negotiation rounds $r$ 3
Max proposal attempts $\text{Att}_{\max}$ 10
Mean-field weight $w_{n,m}$ Inverse hop distance
Consensus threshold $\epsilon$ 0.02 (CPP) / 0 (PC)
Temperature / top-p 0.3 / 1.0
In MACRO-LLM, the rollout horizon $k=2$ balances verification depth against prompt length: the first rollout step performs precise constraint verification on the immediate next state, while the second step applies relaxed feasibility checks to ensure short-term safety (Sec. 3.3). The negotiation rounds $r=3$ are set based on network diameter $D$: since each agent begins with 1-hop neighbor information and each round propagates aggregated statistics by one additional hop, $r$ rounds yield effective coverage of $r+1$ hops. For the PC topologies ($D\leq 4$), $r=3$ provides full network coverage. For the CPP linear topology ($D=7$), the topology-aware partitioning into predecessor and follower subgroups (Theorem 1) reduces the effective propagation distance to $D/2$ per direction, which $r=3$ sufficiently covers.
We evaluate our framework on two complementary domains, CPP and PC, to assess its collaborative performance and generalization capabilities, utilizing high-fidelity simulators for each task. The evaluation covers continuous and discrete actions, linear and graph topologies, fine-grained and long-horizon planning.
G.1 Cooperative Platoon Planning (CPP)
The CPP task is simulated using the Simulation of Urban Mobility (SUMO) [^1] SUMO: https://eclipse.dev/sumo/, licensed under EPL 2.0. (Version 1.21.0).
  • Topology: We model a platoon of eight vehicles ($N=8$) on a single-lane highway segment. The simulation runs for a duration of 60s with 120 time steps in total. Each agent operates with limited perception, accessing only the state information (e.g., headway, velocity) of its immediately adjacent vehicles (predecessor and follower). Due to the linear topology, each agent’s neighborhood naturally partitions into two disjoint subgroups—predecessor and follower—with separate mean-field statistics computed for each direction (Theorem 1).
  • Objective: The control objective is for agents to autonomously adjust acceleration to stabilize the platoon, converging rapidly to an optimal headway of 20 m and a target velocity of 15 m/s. The consensus threshold is set to $\epsilon=0.02$ for the CPP task.
@@ G.2 Pandemic Coordination (PC)
  • Topology: We model the urban environment as a graph of seven agents representing distinct facility types: Home, Office, School, Hospital, Retail, Restaurant, and Government. To construct realistic urban topologies, we utilize Google Maps to identify clusters of functionally similar buildings within specific urban blocks, marking their centroids as individual agents. Communication links are established based on spatial proximity, forming a connected topology as shown in Fig. 6. Since the urban graph does not admit a natural disjoint partition of each agent’s neighborhood, mean-field statistics are computed over the full neighborhood without subgroup decomposition. The simulation spans 120 days, during which agents operate under partial observability regarding global infection rates.
  • Objective: The primary objective is for agents to dynamically adjust regulatory policies (e.g., gathering restrictions, business hours) to minimize the spread of the virus while maintaining essential socio-economic activities. The consensus threshold is set to $\delta=0$ for the PC task.
@@
Appendix H Baseline Methods
To evaluate MACRO-LLM, we implement and adapt the following baselines for both the CPP and PC tasks.
H.1 MARL Baselines
@@ H.2 LLM-based MAS Baselines
We compare against recent frameworks utilizing LLMs for collaboration: (1) ToM-Belief Li et al. (2023b): A MAS that employs second-order Theory of Mind for introspection via belief state updates and environmental feedback; (2) ChatEval Chan et al. (2023): Adopts the Simultaneous-Talk-with-Summarizer protocol (Algorithm 3), which is structurally similar to our task setting; (3) LAMEN Davidson et al. (2024): A framework that constructs internal “mental notes” alongside public messages to facilitate strategic negotiation.
H.3 Adaptation, Fairness, and System Design
To ensure a fair and reproducible comparison, we maintain experimental consistency across all LLM-based methods. In MACRO-LLM, the LLM performs all strategic reasoning: generating collaborative proposals, predicting future observations during rollout verification, evaluating proposal conflicts and updating spatial strategies in the Negotiator, and diagnosing performance drops in the Introspector. Deterministic computations independent of the LLM reasoning process, such as mean-field statistics (Eqs. 23) and drift intensity (Eq. 5), are offloaded to an external numerical module to ensure precision. All inference-dependent computations, including state prediction and constraint evaluation during rollout verification, are performed entirely by the LLM. LLM outputs are validated against a structured XML schema; outputs failing validation trigger an automatic retry (up to 3 attempts).
All LLM-based baselines (ToM-Belief, ChatEval, LAMEN) share the same backbone model (GPT-4o), environmental interfaces, external numerical module, schema validation, and retry mechanism as MACRO-LLM. Methods differ only in their prompting strategies and inter-agent communication protocols (e.g., mental notes generation in LAMEN, belief updates in ToM-Belief). This ensures that performance differences reflect methodological contributions rather than implementation advantages.
Appendix I Case Study: Hierarchical Deployment for Safety-Critical Domains
As a general-purpose strategic coordination framework, MACRO-LLM outputs high-level collaborative directives rather than low-level control commands. For domains with real-time safety constraints, this property allows MACRO-LLM to naturally integrate into a hierarchical deployment paradigm Liu et al. (2025); Wen et al. (2024), where the strategic layer and execution layer operate at different temporal granularities. In this section, we describe the deployment architecture and experimentally validate its feasibility through a decision frequency analysis on the Cooperative Platoon Planning task.
Figure 7: Hierarchical deployment paradigm for MACRO-LLM. The strategic coordination layer executes a full negotiation cycle every $\Delta$ steps to produce collaborative directives, while the domain-specific execution layer operates at every step, applying the cached strategic plan when the state is safe or triggering a safety controller override when a threshold is breached.
I.1 Hierarchical Deployment Paradigm
As illustrated in Fig. 7, the hierarchical paradigm consists of two layers:
  1. 1.
    Strategic Coordination Layer (Second-level granularity). MACRO-LLM operates at a coarser temporal granularity, executing a full negotiation cycle among agents every $\Delta$ time steps. At each decision epoch, it produces strategic directives by reasoning about unobservable states and negotiating with neighboring agents. Between decision epochs, the execution layer carries out the planned actions.
  2. 2.
    Domain-Specific Execution Layer (Millisecond-level granularity). A lightweight controller operates at every time step, performing a safety threshold check on the current state. If the state remains within the safe region, the controller executes the cached strategic plan from the most recent MACRO-LLM decision epoch. If a safety threshold is breached, a domain-specific controller (e.g., PID for CPP) immediately overrides the planned action to ensure system safety.
MACRO-LLM requires no architectural modification to operate within this paradigm—its outputs are inherently strategic directives that can be consumed by any domain-specific execution layer.
I.2 Decision Frequency Analysis
To validate the feasibility of hierarchical deployment, we evaluate MACRO-LLM on the CPP Catch-up scenario under varying decision intervals $\Delta\in\{1,3,5,10\}$, where $\Delta=1$ corresponds to the original setting used in our main experiments. At each decision epoch, MACRO-LLM produces a strategic action spanning the subsequent $\Delta$ time steps. Between decision epochs, the execution layer applies the most recently planned action while monitoring safety constraints via a proportional headway-tracking controller with emergency override. Table 4 presents the results alongside representative baselines from Table 1 for reference.
Table 4: Decision frequency analysis on the CPP Catch-up scenario. DPPO and LAMEN from Table 1 for reference. All metrics $\downarrow$.
Config. Red. RMSE-H RMSE-V SD-H SD-V
DPPO 1.603 0.978 0.994 0.474
LAMEN 1.891 0.648 3.200 0.601
Hier. $\Delta=10$ 90% 1.443 0.623 0.634 0.225
Hier. $\Delta=5$ 80% 1.319 0.721 0.524 0.281
Hier. $\Delta=3$ 67% 1.408 0.596 0.564 0.234
MACRO $\Delta=1$ 0% 1.237 0.721 0.409 0.203
Hier.: hierarchical deployment;
$\Delta$: decision interval (time steps between MACRO-LLM invocations)
Red.: LLM calls reduced relative to $\Delta=1$.
All hierarchical configurations maintain safe operation without any collisions, confirming that the execution layer’s safety override mechanism provides reliable protection between decision epochs. As shown in Table 4, even with 90% fewer LLM invocations ($\Delta=10$), the hierarchical deployment consistently outperforms both DPPO and LAMEN across all metrics. The SD-H of hierarchical deployment ($\Delta=10$) is 36.2% lower than that of DPPO and 80.2% lower than that of LAMEN. This demonstrates that MACRO-LLM’s coordination reasoning provides substantial value even at highly reduced decision frequencies, as the strategic directives continue to guide inter-agent collaboration during the intervals between LLM invocations. Furthermore, increasing the decision frequency yields progressive improvements in coordination quality: SD-H improves from 0.634 ($\Delta=10$) to 0.409 ($\Delta=1$), a 35.5% reduction, confirming that the performance gains stem from MACRO-LLM’s collaborative reasoning. These results indicate that MACRO-LLM can be effectively deployed within a hierarchical paradigm, achieving competitive coordination quality while substantially reducing computational overhead.
Appendix J Detailed Evaluation Metrics
@@ I.2 Decision Frequency Analysis
Appendix K Analysis of Model Generalizability
Table 5: Generalizability analysis of MACRO-LLM across different LLM backbones. S1/S2: Catch-up/Slow-down scenarios of CPP tasks. C1-C3: PC topologies for Helsinki, Hong Kong, and New York.
Metrics ($\downarrow$)
Tasks Base Models RMSE-H RMSE-V SD-H SD-V
CPP S1 DeepSeek-V3.1 1.528 0.553 0.611 0.190
Llama-3.3-70B 1.412 0.779 0.645 0.365
Mistral-3.2-24B 1.891 0.784 0.997 0.326
Qwen3-flash 3.521 1.910 2.935 0.579
GPT-4o-mini 2.906 1.667 1.858 0.936
GPT-4o 1.237 0.721 0.409 0.203
S2 DeepSeek-V3.1 0.649 3.388 0.274 0.417
Llama-3.3-70B 1.163 3.736 0.889 1.224
Mistral-3.2-24B 1.842 3.523 1.157 0.916
Qwen3-flash 3.871 3.678 2.946 0.974
GPT-4o-mini 2.352 3.999 1.598 1.171
GPT-4o 0.496 3.310 0.128 0.238
Tasks Base Models I_n PI_n D_n PD
PC C1 DeepSeek-V3.1 0.098 0.038 0.002 46
Llama-3.3-70B 0.013 0.008 0.000 14
Mistral-3.2-24B 0.043 0.014 0.003 31
Qwen3-flash 0.010 0.008 0.002 16
GPT-4o-mini 0.010 0.010 0.000 17
GPT-4o 0.011 0.008 0.000 10
C2 DeepSeek-V3.1 0.008 0.005 0.001 22
Llama-3.3-70B 0.012 0.005 0.000 16
Mistral-3.2-24B 0.009 0.004 0.001 22
Qwen3-flash 0.008 0.006 0.001 18
GPT-4o-mini 0.006 0.004 0.000 24
GPT-4o 0.008 0.004 0.001 23
C3 DeepSeek-V3.1 0.003 0.003 0.000 11
Llama-3.3-70B 0.013 0.006 0.000 25
Mistral-3.2-24B 0.009 0.006 0.000 14
Qwen3-flash 0.003 0.003 0.000 13
GPT-4o-mini 0.003 0.003 0.000 13
GPT-4o 0.004 0.003 0.000 25
We investigate the robustness of MACRO-LLM by instantiating it with diverse LLM backbones, including DeepSeek-V3.1 Liu et al. (2024), Llama-3.3-70B-Instruct Grattafiori et al. (2024), Mistral-Small-3.2-24B-Instruct, Qwen3-flash Bai et al. (2023), GPT-4o-mini, and GPT-4o. The results are detailed in Table 5. In the CPP task, performance gaps across backbones are primarily attributable to the fine-grained continuous control demands: GPT-4o and DeepSeek-V3.1 reduce RMSE-H by over 47.4% compared to GPT-4o-mini in the Catch-up scenario. Conversely, in the PC task, Qwen3-flash and GPT-4o-mini achieve containment efficacy comparable to GPT-4o across all three topologies, as this task relies more on strategic planning than on numerical precision. This contrast confirms that the observed performance variation stems from task-specific precision requirements rather than framework limitations, and that MACRO-LLM’s structured negotiation mechanism enables even lightweight models to accomplish complex multi-agent coordination tasks.
Appendix L Computational and Communication Scalability Analysis
@@
We provide a theoretical and experimental analysis of MACRO-LLM, focusing on communication complexity and computational workload.
L.1 Communication Complexity Analysis
@@
Following the definition by Yao Yao (1979), communication complexity is quantified by the total number of bits exchanged during the protocol. In our decentralized framework, we focus on the per-agent communication load. Let $\mathcal{G}=(\mathcal{V},\mathcal{E})$ be the agent communication graph. We denote the average node degree as $\bar{d}$ and the maximum negotiation rounds as $r$. The topology of the CPP task is linear ($\bar{d}\approx 2$). In the PC task, the topology represents an urban graph with average degree $\bar{d}_{\text{PC}}$. We assume sparse topologies where $\bar{d}$ does not grow linearly with the total number of agents $N$ (i.e., $\bar{d}\ll N$). As the network scales up, MACRO-LLM employs mean-field approximation to compress the states of unobservable agents into a single virtual node representation. Neglecting inference stochasticity, the per-agent message size is theoretically invariant to the network scale, which is represented as $M$. Consequently, the per-agent communication complexity can be represented as $O(\bar{d}\cdot r\cdot M)$. Since $\bar{d}$, $r$, and $M$ are independent of network scale $N$, the per-agent load remains constant as the network scales in theory. In other word, the per-agent communication load does not increase with the number of agents as long as the topology’s connectivity degree remains constant. The scalability results in Table 6 demonstrate as the platoon scales from $N=8$ to $N=32$, coordination metrics remain stable with no monotonic degradation.
L.2 Computational Efficiency and Latency
Given that the input prompt length remains bounded as analyzed above, the upper bound of the inference workload for a single agent does not increase with the number of agents. Under API-based deployment, the per-step latency is dominated by network round-trip overhead of remote API calls rather than the framework’s reasoning complexity itself. The per-step latency is dominated by API network overhead of remote LLM calls ($\sim$5.99s per call for GPT-4o). This overhead can be substantially mitigated through two orthogonal and composable strategies:
(1) Local deployment of lightweight models. Table 5 demonstrates that lightweight models such as GPT-4o-mini and Qwen3-flash achieve competitive coordination performance across both tasks. Locally deploying these models eliminates API network latency entirely, reducing per-call inference time by an order of magnitude compared to remote API access.
(2) Hierarchical deployment with reduced decision frequency. As demonstrated in Appendix I, MACRO-LLM maintains superior performance over all baselines even when decision frequency is reduced by 90% (Table 4), confirming that the framework’s coordination value does not depend on high-frequency invocation.
These two strategies are composable: combining local lightweight models with increased decision intervals can reduce the effective latency by over an order of magnitude relative to the default API-based configuration.
Furthermore, MACRO-LLM exhibits instant zero-shot adaptation to new scenarios, whereas MARL baselines necessitate extensive retraining when transferring to unseen topologies. For dynamic, open-world deployments where frequent retraining is impractical, this zero-shot capability represents a significant practical advantage that offsets the per-step inference cost.
L.3 Scalability Metrics Across Fleet Sizes
Fig. 3 represents the coordination performance for the CPP Catch-up scenario as the fleet size increases from $N=8$ to $N=32$. Table 6 reports the detailed metrics.
Table 6: Scalability analysis on CPP Catch-up with GPT-4o.
@@
8 1.237 0.721 0.409 0.203
16 1.003 0.749 0.493 0.355
24 0.908 0.664 0.520 0.382
32 0.728 0.628 0.433 0.397
All metrics remain stable as $N$ increase to 32, with no monotonic degradation. The results indicate that the mean-field aggregation mechanism effectively compresses large-scale interactions without sacrificing coordination precision.
Appendix M Extended Ablation Analysis
@@
M.1 Component Analysis in CPP Task
The removal of specific modules leads to critical performance degradation in the high-precision and highly dynamic CPP task:
@@
M.2 Component Analysis in PC Task
A similar pattern holds for the PC task:
@@
Appendix N Analysis of Robustness
Table 7: Robustness analysis of MACRO-LLM with GPT-4o on the CPP task (5 independent runs). Results are reported as mean $\pm$ standard deviation.
Scenario RMSE-H ($\downarrow$) RMSE-V ($\downarrow$) SD-H ($\downarrow$) SD-V ($\downarrow$)
Catch-up 1.237 $\pm$ 0.068 0.721 $\pm$ 0.124 0.409 $\pm$ 0.034 0.203 $\pm$ 0.054
Slow-down 0.496 $\pm$ 0.044 3.310 $\pm$ 0.010 0.128 $\pm$ 0.054 0.238 $\pm$ 0.073
Table 8: Robustness analysis of MACRO-LLM with GPT-4o on the PC task (5 independent runs). Results are reported as mean $\pm$ standard deviation.
City I_n ($\downarrow$) PI_n ($\downarrow$) D_n ($\downarrow$) PD_n* ($\downarrow$)
Helsinki 0.011 $\pm$ 0.001 0.008 $\pm$ 0.000 0.000 $\pm$ 0.000 0.081 $\pm$ 0.024
Hong Kong 0.008 $\pm$ 0.004 0.004 $\pm$ 0.001 0.001 $\pm$ 0.001 0.190 $\pm$ 0.063
New York 0.004 $\pm$ 0.001 0.003 $\pm$ 0.000 0.000 $\pm$ 0.000 0.204 $\pm$ 0.117
* PD_n = PD / total simulation days, representing the proportion of pandemic duration relative
to the total evaluation period.
To evaluate the stability of MACRO-LLM, we conduct 5 independent runs with GPT-4o and report the mean and standard deviation in Tables 7 and 8. For the CPP task, MACRO-LLM exhibits low variance across all metrics. In the Catch-up scenario, the SD-H of $0.409\pm 0.034$ is 58.9% lower than that of DPPO, and even the upper bound ($0.443$) remains 55.4% lower, confirming that the stability advantage in Table 1 is not attributable to favorable randomness. Similarly, in the Slow-down scenario, the RMSE-H ($0.496\pm 0.044$) is 46.3% lower than that of LAMEN and 71.2% lower than that of DPPO across all runs. For the PC task, infection-related metrics (I_n, PI_n, D_n) remain tightly concentrated across all three topologies, with standard deviations at or near zero. While PD_n shows relatively higher variance in more complex topologies (e.g., New York: $0.204\pm 0.117$), this is expected given the larger state space and stochastic disease dynamics inherent to denser urban networks. These results confirm that MACRO-LLM achieves consistently superior performance with reliable reproducibility across independent trials.
Appendix O Analysis of Failure Cases and Solutions